OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2

OpenStudy (amorfide):

I think you mean y-y1=m(x-x1) if so then you will replace y1 with the y coordinate replace m with the slope value and replace x1 with x coordinate

OpenStudy (anonymous):

the choices are: y − 6 = (x + 3) y = x − 6 y + 3 = (x − 6) x − 2y = 12

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

i gave you the equation y-y1=m(x-x1) sub in y1 which is the value of the y coordinate sub in m which is the value of the slope sub in x1 which is the x coordinate then simplify

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

I got b is that right?

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

i dont get any of those answers lol but sure go with b, it's the closest i guess

OpenStudy (andriod09):

point/slope is this equation: \[\frac{\Delta y}{\Delta x}\]

OpenStudy (andriod09):

so then the equation filled in would be: \[\frac{\Delta(-3)}{\Delta6}\]

OpenStudy (anonymous):

so b?

OpenStudy (andriod09):

does that fit the equation: \[y=mx+b\] that should help you, and yes it does.

OpenStudy (anonymous):

so it is or is not b?

OpenStudy (amorfide):

x-2y=12

OpenStudy (anonymous):

are you sure? nvmind i see how you got that okay thanks yall